OpenStudy (anonymous):

Simplify 4x/9 + x-10/9 and tell me how !

OpenStudy (anonymous):

how did you get that ?

OpenStudy (anonymous):

by adding you equation boyyyy!

OpenStudy (anonymous):

the fink looked at the wrong equation god darn!!

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[\frac{ 4x }{ 9 } +\frac{ (x- 10) }{ 9 }\] so you will have to add \[\frac{ 5x - 10 }{ 9 }\] then it simplifies to \[\frac{ 5(x -2) }{ 9 }\]

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

alright thank you

OpenStudy (anonymous):

yw